OpenStudy (anonymous):

Find the slope of the line that passes through the points (0, 0) and (-2, -3). 2/3 3/2 -3/2

OpenStudy (anonymous):

Slope equation= (y2-y1)/(x2-x1)

OpenStudy (anonymous):

slope equation = (x2-x1)(y2-y1)=r2